What constant force is needed in order to increase the speed of a car from 30 km/h to 

60 km/h in a time of 6s if the mass of the car is 2400 kg?



1
Expert's answer
2021-07-05T15:23:43-0400
"F=ma,""F=\dfrac{m(v-v_0)}{t},""F=\dfrac{2400\ kg\cdot(16.67\ \dfrac{m}{s}-8.33\ \dfrac{m}{s})}{6\ s}=3336\ N."